OpenStudy (anonymous):

Use addition or substitution to find the value of y for this set of equations. 2x - 3y = -15 x = 4y

OpenStudy (anonymous):

you could use either but substitution would be easier

OpenStudy (anonymous):

use x=4y to substitute in the equation 2x-3y=-15 which will be: 2(4y)-3y=-15 8y-3y=-15 5y=-15 y=-3 ifyou substitute y=-3 in the original equation you will get the value of x which is -12 Good luck